### v1

@@ Abstract.
The rapid proliferation of large language models (LLMs) as general purpose chatbots available to the public raises hopes around expanding access to professional guidance in law, medicine, and finance, while triggering concerns about public reliance on LLMs for high-stakes circumstances. Prior research has speculated on high-level ethical considerations but lacks concrete criteria determining when and why LLM chatbots should or should not provide professional assistance. Through examining the legal domain, we contribute a structured expert analysis to uncover nuanced policy considerations around using LLMs for professional advice, using methods inspired by case-based reasoning. We convened workshops with 20 legal experts and elicited dimensions on appropriate AI assistance for sample user queries (“cases”). We categorized our expert dimensions into: (1) user attributes, (2) query characteristics, (3) AI capabilities, and (4) impacts. Beyond known issues like hallucinations, experts revealed novel legal problems, including that users’ conversations with LLMs are not protected by attorney-client confidentiality or bound to professional ethics that guard against conflicted counsel or poor quality advice. This accountability deficit led participants to advocate for AI systems to help users polish their legal questions and relevant facts, rather than recommend specific actions. More generally, we highlight the potential of case-based expert deliberation as a method of responsibly translating professional integrity and domain knowledge into design requirements to inform appropriate AI behavior when generating advice in professional domains.
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
@@ 1. Introduction
As large language models (LLMs) rapidly evolve, their capabilities are increasingly integrated into complex professional domains like the legal sector. LLM-based chatbots offering legal advice have emerged as a potential tool to improve accessibility and personalize legal services (Guo et al., 2023; Schmitz and Zeleznikow, 2022). However, the reliance on imperfect AI models for high-stakes decisions introduces significant risks, such as misleading users with inaccurate information, security breaches, and biases in advice (Wendel, 2019; Yue et al., 2023). In this regard, the EU AI Act officially designates AI systems used for “assistance in legal interpretation and application of the law” as high-risk (European Commission, 2021).
@@
However, most prior research in this field speculates on high-level concerns such as inaccuracy and real-world impacts (Zheng et al., 2021; Lin et al., 2021; Valvoda et al., 2023; Ling et al., 2023; Ray, 2023) without elucidating concrete criteria for when and why AI should or should not provide professional advice to users. This fails to produce actionable design requirements that can meaningfully guide real-world AI deployment practices surrounding professional advice systems. Borrowing domain knowledge and drawing voices from direct stakeholders to inform AI responsible governance is an emergent practice (Antoniak et al., 2023) that has not yet been applied to the legal realm. We aim to address this gap by directly drawing from the voices of legal professionals to translate domain knowledge into actionable AI deployment guidance. Our research questions are:
@@
  • RQ1: What key considerations do legal professionals identify in determining appropriate AI responses to lay users’ legal questions?
@@
  • RQ2: What guiding principles and response strategies do legal professionals recommend for AI systems providing legal advice to lay users?
@@
Figure 1. Overview of our research process and findings. We collected 33 "cases," meaning realistic user queries, and generated 7 response strategies. During workshops, 20 experts provided their opinions on appropriate AI response strategies and the key dimensions they considered for their judgments. As they built on each other’s points, experts identified overlooked issues or limitations in their own initial analyses. We conducted qualitative and quantitative analyses of pre-survey results, workshop transcripts, and workshop documents. Grounding our findings in literature across LLMs, law, and AI ethics/policy, we developed a clear 4-dimensional framework that informed expert judgment and provided guiding principles for appropriate AI legal advice.
@@
We leverage a process (Figure 1) inspired by case-based reasoning, an approach commonly used in pedagogical material for a wide variety of fields, including law and moral theory (Chen and Zhang, 2023; Feng et al., 2023; Kolodner, 1992; Paulo, 2015; Jonsen, 1986), to enable discussion of ethical considerations grounded in concrete cases. We convened 7 interactive workshops with 20 legal experts by providing them with 33 queries (“cases”) and sampled from a wide variety of LLM chatbot response strategies, from outright refusal to respond to concrete opinions regarding specific actions to take. Through analysis of the collected data, iterative rounds of discussion among authors, and literature review across the fields of law, natural language processing (NLP), and AI ethics, we consolidated and identified the significant dimensions that guided experts’ evaluations and guiding principles for desirable AI responses.
@@
For RQ1, we identified 25 key dimensions that should inform potential AI responses (Figure 3). We classified dimensions into four categories—(1) user attributes and behaviors, (2) query characteristics, (3) current AI capabilities, and (4) impact considerations. For RQ2, experts expressed their preferences on a spectrum of sampled responses spanning from blanket refusals to detailed and actionable opinions. Key tensions emerged around offering the possibility of personalized advice beyond factual legal information and engaging multi-turn dialogue through follow-up questions to polish users’ questions and distill relevant facts. Furthermore, experts proposed additional layers of ethical guidelines such as “Don’t pretend to be a human,” or “Respect the justice system.”
@@
Our contributions are multifold: First, we demonstrate how our case-based expert deliberation process was effective in leveraging experts’ knowledge and experience to elicit a rich set of dimensions. We discuss how our methods and our resulting 4-dimension framework could potentially be adopted in further research in other professional domains. Second, our 4-dimension framework, spanning across query-specific concerns to more systemic constraints grounded in legal and technical literature, provides a fertile groundwork for AI policy creation beyond speculative theoretical principles. Third, in addition to dimensions, we portray expert disagreements on appropriate AI responses, while highlighting where experts agreed on information-focused or multi-turn issue-spotting approaches. Finally, we reveal novel legal and ethical considerations, such as unauthorized practice of law, confidentiality, and liability for inaccurate advice, overlooked in the LLM ethics literature. This illustrates that responsible AI legal advice requires a cross-disciplinary synthesis that spans technology, law, and ethics, learning from accumulated knowledge in professional communities.
@@ 2. Related Work and Our Approach
To produce responsible LLMs for legal advice, we must examine both the capabilities and limitations of state-of-the-art LLM technology and legal services historically regulated by stringent professional ethics rules. Hence, our research is indebted to a substantial body of scholarship in the fields of NLP, law, and AI ethics/policy.
@@
LLMs’ Promises and Limitations. LLMs have shown immense promise in lowering historic barriers to services that have long relied on highly-trained human specialists (Metzler et al., 2021), across domains like healthcare (Antoniak et al., 2023; Snoswell et al., 2023; Sallam, 2023; Javaid et al., 2023; Singhal et al., 2023), finance (Northey et al., 2022; Suhel et al., 2020), and law (Greco and Tagarelli, 2023; Valvoda et al., 2023; Nay, 2023). Our focus is on law, a field where human attorneys undergo years of education to provide counsel largely out of reach to lay people (Stockdale and Mitchell, 2019; Wendel, 2019). Researchers have endeavored to enhance legal prediction and reasoning through dedicated datasets, in-domain fine-tuning, and prompt engineering (Greco and Tagarelli, 2023; Liga and Robaldo, 2023; Nay et al., 2023; Huang et al., 2023; Valvoda et al., 2023; Nguyen et al., 2023; Guha et al., 2023; Prasad et al., 2022; Trautmann et al., 2022). However, ensuring accuracy and high-quality writing remains a challenge (Trozze et al., 2023, 2023). Most critically, as statistical models, LLMs can “hallucinate” answers not grounded in their training data, severely compromising reliability (Shen et al., 2023; Liu et al., 2023; Metzler et al., 2021). Furthermore, researchers stress the lack of security (Iqbal et al., 2023; Yao et al., 2023) and interpretability (Singh et al., 2023; Saha et al., 2023; Zhao et al., 2023), alongside issues of bias and stereotypes in language models and their datasets (Ghosh and Caliskan, 2023; Kotek et al., 2023; Omrani Sabbaghi et al., 2023; Dhingra et al., 2023). While advances have been made, rigorous examination of risks is needed given that flawed legal counseling can severely infringe on rights, livelihoods, and liberties (Ling et al., 2023; Ray, 2023). We elicit these risks and other key dimensions for determining the appropriateness of AI responses directly from legal experts in our work.
@@
Legal Doctrines Governing Legal Advice. If AI systems attain human-level accuracy, could AI’s legal advice be provided to users without any legal concerns? Law scholars have intensely debated these issues, since much before the rise of LLMs, when people imagined AI judges and attorneys (Sourdin, 2018; Simshaw, 2018; Medianik, 2017). The most common doctrines involved are unauthorized practice of law (UPL) and professional ethics rules (Akata et al., 2020; Sil et al., 2019; Lightbourne, 2017; Medianik, 2017; Lightbourne, 2017). The states governing attorneys’ licenses prohibit unlicensed individuals from providing legal advice to others (cal, 1970). For instance, California law allows paralegals to do fact-gatehring and retreiving “information,” but not to provide “legal advice.” (cal, 2007) Applying this rule to AI systems, Spahn argues non-lawyers using AI to provide legal advice or prepare documents for third parties could violate the UPL (Spahn, 2017), while Stockdale & Mitchell found legal advice privilege may still apply between users and AI chatbots in some jurisdictions (Stockdale and Mitchell, 2019). Moreover, reflecting on professional ethics, Haupt stresses AI’s professional advice must demonstrate competence, trust, responsibility, and ethics (Haupt, 2019). Our work extends these discussions to contemporary AI systems powered by LLMs.
@@
Responsible AI Ethics/Policy. Researchers have endeavored to propose “guardrails” to prevent the unethical or unjust outcome caused by LLMs. Much of the pioneering work categorizes key challenges such as inaccuracy, bias in models, inequality, over-reliance, and explainability (Bender et al., 2021; Ray, 2023; D’Arcy and Bender, 2023; Shah and Bender, 2022; Ling et al., 2023; Solaiman et al., 2023). Some work extends to clarifying specific guidelines such as Singha & Bender (Shah and Bender, 2022) (e.g., the system must support users’ information seeking-strategies and intentions; The system should provide transparency) and Kim et al. (Kim et al., 2023b) (e.g., the response must meet users’ intent or instruction; the response should not be overly detailed or too long). Antoniak et al. (Antoniak et al., 2023) outline guiding principles for NLP in healthcare (e.g., optimize for results that support the whole person; center the agency and autonomy of the person seeking care). Our work aligns with Antoniak et al.’s approach by first diving into domain-specific concerns (legal advice) associated with LLM-generated responses. We then identify aspects of our findings that are applicable to and enrich the discussion of more guardrails for AI.
@@
Eliciting Expert Knowledge and Case-based Reasoning. Incorporating experts’ domain knowledge into AI development has emerged as a critical “participatory AI” approach (Balaram et al., 2018; Costanza-Chock, 2018; Durmus et al., 2023; Birhane et al., 2022; Delgado et al., 2023; Queerinai et al., 2023; Nekoto et al., 2020). Researchers have facilitated expert discussions to evaluate sociotechnical implications of LLMs (Singhal et al., 2023; Solaiman et al., 2023; Peskoff and Stewart, 2023; Antoniak et al., 2023). However, unlike prior work focusing on theoretical ethical principles (Solaiman et al., 2023; Antoniak et al., 2023; Ganguli et al., 2022) or post-hoc system evaluation (Veen et al., 2023; Peskoff and Stewart, 2023; Balas et al., 2023), we pursued the case-based approach to spurring expert deliberation based on their clinical experience. We present legal professionals with realistic legal queries that an AI systems could receive from lay end-users. We thus take a case-based reasoning approach (Chen and Zhang, 2023; Feng et al., 2023), informed by moral philosophy (Kolodner, 1992; Mackie, 2003; Paulo, 2015; Jonsen, 1986; Smith, 1976; Fullinwider, 2010) and legal theories (Cardozo and Kaufman, 2010; Grey, 1983; Sunstein, 2018) that emphasizes case-by-case judgments to shape guidelines instead of universally applying top-down rules. Distinguished from most AI policy guidelines provide a single set of universally-agreeable principles (Antoniak et al., 2023), this approach enables us to highlight critical value-laden topics that experts disagreed with each other, while proposing unique dimension framework, ranging from case-specific concerns to structural constraints, which experts considered to determine proper AI responses.
@@ 3. Methods: Case-based Expert Deliberation
We conducted seven small-group workshops via Zoom with 20 expert participants in August 2023. We assumed a scenario involving general-purpose conversational AI systems like ChatGPT or Bing Chat available to lay users, different from professional AI tools assisting legal practitioners.
@@
Recruitment. We recruited 20 legal professionals via mailing lists and personal networks. Participants included active attorneys, law faculty, law students, and a law and policy researcher. Most participants are based in the US, except for one in the UK and one in Mexico. The cohort spans early-career to lawyers over 20 years of experience, with varying degrees of general and professional AI usage. Table 1 summarizes participants’ backgrounds and self-reported AI usage patterns. More detailed information is available at Appendix B.
@@
Construction of Cases. We manually sourced 33 cases from a combination of (1) the popular subreddit r/legaladvice (with wording edited slightly for anonymization and clarity), and (2) existing cases in case studies and case law familiar to our team member who is a practicing attorney. Our cases covered facets of law most relevant to lay users, including criminal law, employment disputes, housing issues, corporate governance, and family law. This diversity of user intents (e.g., getting out of trouble, advocating for others, minimizing their costs), impacted third parties (employers, colleagues, landlords, family members, protesters), and the degree of damage (physical, financial, mental), intended to elicit a wide range of discussion across legally and ethically sensitive contexts. Our cases can be viewed at https://anonymous.4open.science/r/case-law-ai-policy-4CB5/data/cases.csv.
@@
Workshop Procedures. During the workshop, we presented 20 randomly-chosen cases along with 7 generic response strategies for AI responses on a shared Google document. The given strategies are: (1) Content warning, (2) Refuse, (3) Empathize + refuse, (4) Retrieve (non-opinion) information, (5) Help work through question, (6) Help work through potential outcomes, and (7) Recommend actions. We provided an example response for each strategy, and examples were empirically derived from what we observed from OpenAI’s GPT-3, 3.5, and 4. Because GPT tends to refuse to give detailed advice such as options (6) and (7), we drafted more specific answers complemented by known prompt engineering techniques such as drawing a hypothetical scenario (Abdelnabi et al., 2023). Response strategies and corresponding examples are available in Appendix A. Figure 2 provides an overview of the Google document we gave to the participants. After an introduction, each participant was given 10–15 minutes to freely choose 2–4 cases and (1) select the proper AI response strategies or produce their own preferred response and (2) the key dimensions impacting their decision in an individual workspace. Then, experts had 30–35 minutes to discuss with each other why they chose certain response strategies and what dimensions they took into account to determine the proper strategies.
@@
Figure 2. Overview of case examples and AI response strategies and examples provided to participants. Participants were given 10–15 minutes to review  20 legal case prompts on a shared document, select 2–4 cases to examine further, and specify appropriate AI responses and influential considerations in their individual workspace on the same document.
@@
Analysis. We analyzed Google Docs and transcripts using abductive coding (Tavory and Timmermans, 2014). Integrating both empirical data and available theory in an iterative process, our findings are informed by and enter into dialogue with literature in law, NLP, and AI ethics/policy, synthesizing relevant aspects of these fields within the context of our research questions. Based on their analysis of 2 transcripts, two authors initially developed a codebook of dimensions and responses, informed by the human-AI-context framework (Kim et al., 2023a), legal professional ethics literature (e.g., Haupt, 2019; Wendel, 2019), and ethical concerns in LLM interactions (e.g., Kim et al., 2023b; Valvoda et al., 2023). Following the codebook finalized through multiple all-author meetings, two coders independently analyzed the data and meticulously cross-checked each other’s work. This process, where both coders examined all documents and reached consensus on coding, rendered inter-rater reliability metrics unnecessary (McDonald et al., 2019).

### v2

@@ Abstract.
Large language models (LLMs) are increasingly capable of providing users with advice in a wide range of professional domains, including legal advice. However, relying on LLMs for legal queries raises concerns due to the significant expertise required and the potential real-world consequences of the advice. To explore when and why LLMs should or should not provide advice to users, we conducted workshops with 20 legal experts using methods inspired by case-based reasoning. The provided realistic queries (“cases”) allowed experts to examine granular, situation-specific concerns and overarching technical and legal constraints, producing a concrete set of contextual considerations for LLM developers. By synthesizing the factors that impacted LLM response appropriateness, we present a 4-dimension framework: (1) User attributes and behaviors, (2) Nature of queries, (3) AI capabilities, and (4) Social impacts. We share experts’ recommendations for LLM response strategies, which center around helping users identify ‘right questions to ask’ and relevant information rather than providing definitive legal judgments. Our findings reveal novel legal considerations, such as unauthorized practice of law, confidentiality, and liability for inaccurate advice, that have been overlooked in the literature. The case-based deliberation method enabled us to elicit fine-grained, practice-informed insights that surpass those from de-contextualized surveys or speculative principles. These findings underscore the applicability of our method for translating domain-specific professional knowledge and practices into policies that can guide LLM behavior in a more responsible direction.
journalyear: 2024copyright: rightsretainedconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3659048isbn: 979-8-4007-0450-5/24/06
@@ 1. Introduction
Human-like conversational capabilities and the vast knowledge of large language models (LLMs) have shown promise in improving access to services traditionally requiring human specialists (Metzler et al., 2021), in domains such as healthcare (Antoniak et al., 2023; Snoswell et al., 2023; Sallam, 2023; Javaid et al., 2023; Singhal et al., 2023), finance (Northey et al., 2022; Suhel et al., 2020), and law (Greco and Tagarelli, 2023; Valvoda et al., 2023; Nay, 2023). In the legal field, where attorneys undergo extensive training to provide counsel, often beyond the reach of lay people (Stockdale and Mitchell, 2019; Wendel, 2019), LLM-based chatbots offering legal advice have emerged as a potential accessibility aid (Guo et al., 2023; Schmitz and Zeleznikow, 2022). However, relying on imperfect LLMs for high-stakes legal decisions raises concerns around low-quality advice and privacy risks (Wendel, 2019; Yue et al., 2023). These concerns have prompted the EU AI Act to designate AI systems used for “assistance in legal interpretation and application of the law” as “high-risk” (European Commission, 2021).
@@
Most prior research in this field speculates on high-level concerns such as inaccuracy and real-world impacts (Zheng et al., 2021; Lin et al., 2021; Valvoda et al., 2023; Ling et al., 2023; Nazir and Wang, 2023). However, these studies rarely articulate concrete criteria for when and why LLMs should or should not provide professional advice to users. As a result, they offer insufficient guidance to produce actionable design requirements that can inform real-world LLM deployment practices. Incorporating perspectives from domain experts is an emerging approach for responsible LLM policies (Antoniak et al., 2023), but has not been applied to the legal domain. Our work aims to bridge this gap by addressing the following questions:
@@
  • RQ1: What key dimensions do legal professionals identify in determining appropriate LLM responses to lay users’ legal questions?
@@
  • RQ2: What guiding principles and response strategies do legal professionals recommend for LLM systems providing legal advice to lay users?
@@
Figure 1. Overview of our research process and findings. We collected 33 “cases,” meaning realistic user queries, and provided 7 response strategies. During workshops, 20 experts provided their opinions on appropriate LLM response strategies and the key dimensions they considered for their judgments. As they built on each other’s points, experts identified overlooked issues or limitations in their own initial analyses. We qualitatively and quantitatively analyzed workshop data and pre-survey results. Grounding our findings in literature across LLMs, law, and AI ethics/policy, we developed a clear 4-dimension framework that informed expert judgment and provided guiding principles for appropriate LLM legal advice.
@@
We leverage a process (Figure 1) inspired by case-based reasoning, an approach commonly used in pedagogical material for a wide variety of fields, including law and moral theory (Chen and Zhang, 2023; Feng et al., 2023; Kolodner, 1992; Paulo, 2015; Jonsen, 1986), to enable discussion of ethical considerations grounded in concrete cases. We convened 7 interactive workshops with 20 legal experts by providing them with 33 queries (“cases”) and asked them to evaluate 7 simulated responses that could arise from LLM chatbots, ranging from outright refusal to recommendation of specific actions with legal judgment. Through analysis of the collected data, iterative rounds of discussion among authors, and literature review across the fields of law, natural language processing (NLP), and AI ethics, we consolidated and identified the significant dimensions that affected experts’ evaluations and guiding principles for desirable LLM responses.
@@
For RQ1, we identified 25 key dimensions that should inform potential LLM responses (Figure 3). We classified dimensions into four categories—(1) User attributes and behaviors, (2) Nature of queries, (3) AI capabilities, and (4) Social impacts. For RQ2, experts generally expressed their preferences for information-focused responses. Instead of seeking definitive legal judgement, some suggested leveraging LLMs’ multi-turn dialogue capabilities to polish users’ questions and distill relevant facts through follow-up questions. Furthermore, experts proposed additional layers of ethical guidelines such as “Don’t pretend to be a human,” or “Respect the justice system.”
@@
Our contributions are multi-fold: First, our 4-dimension framework, spanning across query-specific concerns to more systemic constraints grounded in legal and technical literature, provides a fertile groundwork for LLM policy creation beyond speculative theoretical principles. Second, in addition to dimensions, we portray expert disagreements on appropriate LLM responses, while highlighting where experts agreed on information-focused or multi-turn issue-spotting approaches. Third, we demonstrate how our case-based expert deliberation process was effective in leveraging experts’ knowledge and experience to elicit a rich set of dimensions. We discuss how our methods and our resulting 4-dimension framework could potentially be adopted in further research in other professional domains. Finally, we reveal novel legal and ethical considerations, such as unauthorized practice of law, confidentiality, and liability for inaccurate advice, overlooked in the LLM literature. This illustrates that responsible AI legal advice requires a cross-disciplinary synthesis that spans technology, law, and ethics, learning from accumulated knowledge in professional communities.
@@ 2. Related Work and Our Approach
To develop policies for LLMs providing legal advice, we must consider both the current capabilities and limitations of LLM technology, as well as existing legal ethics rules aimed at preventing harms from improper legal advice. Our research builds upon prior work in the fields of NLP, law, and AI ethics and policy.
@@
LLMs’ Promises and Limitations. Researchers have endeavored to enhance legal prediction and reasoning through dedicated datasets, in-domain fine-tuning, and prompt engineering (Greco and Tagarelli, 2023; Liga and Robaldo, 2023; Nay et al., 2023; Huang et al., 2023; Valvoda et al., 2023; Nguyen et al., 2023; Guha et al., 2023; Prasad et al., 2022; Trautmann et al., 2022). However, ensuring accuracy and high-quality writing remains a challenge (Trozze et al., 2024, 2024). Most critically, as statistical models, LLMs can “hallucinate” answers not grounded in their training data, severely compromising reliability (Shen et al., 2023; Liu et al., 2023; Metzler et al., 2021). Furthermore, researchers stress the lack of security (Iqbal et al., 2023; Yao et al., 2023) and interpretability (Singh et al., 2023; Saha et al., 2023; Zhao et al., 2024), alongside issues of bias and stereotypes in their datasets (Ghosh and Caliskan, 2023; Kotek et al., 2023; Omrani Sabbaghi et al., 2023; Dhingra et al., 2023). While advances have been made, rigorous examination of risks is needed given that flawed legal counseling can severely infringe on rights, livelihoods, and liberties (Ling et al., 2023; Trozze et al., 2024; Alarie and McCreight, 2023). Our work contributes to this critical assessment by eliciting insights from legal experts on these risks and other key dimensions for determining the appropriateness of LLM responses in legal contexts.
@@
Doctrines Governing Legal Advice. The provision of legal advice by AI systems raises significant questions related to legal doctrines. Law scholars have intensely debated these issues, since much before the rise of LLMs, when people imagined AI judges and attorneys (Sourdin, 2018; Simshaw, 2018; Medianik, 2017). The most common doctrines involved are unauthorized practice of law (UPL) and professional ethics rules (Akata et al., 2020; Sil et al., 2019; Lightbourne, 2017; Medianik, 2017). The states prohibit unlicensed individuals from providing legal advice to others (cal, 1970). For instance, California law allows paralegals to do fact-gathering and retrieving “information,” but not to provide “legal advice.” (cal, 2007) Applying this rule to AI systems, Spahn argues non-lawyers using AI to provide legal advice or prepare documents for third parties could violate the UPL (Spahn, 2017), while Stockdale & Mitchell finds that legal advice privilege may still apply between users and AI chatbots in some jurisdictions (Stockdale and Mitchell, 2019). Reflecting on professional ethics, Haupt stresses AI’s professional advice must demonstrate competence, trust, responsibility, and ethics (Haupt, 2019). Our work extends these discussions to state-of-art conversational LLM systems.
@@
AI Ethics/Policy. Researchers have endeavored to propose “guardrails” to prevent the unethical or unjust outcomes caused by LLMs. Much of the pioneering work categorizes key challenges such as inaccuracy, bias in models, inequality, over-reliance, and explainability (Bender et al., 2021; Nazir and Wang, 2023; D’Arcy and Bender, 2023; Shah and Bender, 2022; Ling et al., 2023; Solaiman et al., 2023). Some work extends to clarifying specific guidelines such as Shah & Bender (Shah and Bender, 2022) (e.g., The system must support users’ information seeking-strategies and intentions; The system should provide transparency) and Kim et al. (Kim et al., 2023a) (e.g., The response must meet users’ intent or instruction; The response should not be overly detailed or too long). Antoniak et al. (Antoniak et al., 2023) outline guiding principles for NLP in healthcare (e.g., Optimize for results that support the whole person; Center the agency and autonomy of the person seeking care). Our work also aims to produce actionable principles that guide LLM behavior reflecting on domain-specific concerns in legal advice.
@@
Eliciting Expert Knowledge and Case-based Reasoning. Incorporating experts’ domain knowledge into AI development has emerged as one of the “participatory AI” approaches (Balaram et al., 2018; Costanza-Chock, 2020; Durmus et al., 2023; Birhane et al., 2022; Delgado et al., 2023; Queerinai et al., 2023; Nekoto et al., 2020). Researchers have facilitated expert discussions to evaluate the sociotechnical implications of LLMs (Singhal et al., 2023; Solaiman et al., 2023; Peskoff and Stewart, 2023; Antoniak et al., 2023). Unlike prior work focusing on high-level ethical principles (Solaiman et al., 2023; Antoniak et al., 2023; Ganguli et al., 2022) or post-hoc system evaluation (Veen et al., 2023; Peskoff and Stewart, 2023; Balas et al., 2023), we pursued the case-based reasoning (Chen and Zhang, 2023; Feng et al., 2023) approach to spur expert deliberation based on their clinical experience. We present legal professionals with realistic legal queries that LLM systems could receive from lay end-users. This approach comes from moral philosophy (Kolodner, 1992; Mackie, 2003; Paulo, 2015; Jonsen, 1986; Smith, 1982; Fullinwider, 2010) and legal theories (Cardozo and Kaufman, 2010; Grey, 1983; Sunstein, 2018) that emphasizes case-by-case judgments to shape guidelines instead of applying top-down rules. Distinguished from most AI policy guidelines that provide a single set of universally-agreeable principles (Antoniak et al., 2023), case-based deliberation enables us to highlight critical value-laden topics on which experts disagreed with each other. Furthermore, it allows us to synthesize a dimensional framework, ranging from case-specific concerns to structural constraints, which experts considered to determine proper LLM responses.
@@ 3. Methods: Case-based Expert Deliberation
We conducted seven small-group workshops on Zoom with 20 expert participants in August 2023. We assumed a scenario involving general-purpose conversational LLM systems like ChatGPT or Bing Chat available to lay users, different from professional tools assisting legal practitioners.
@@
Recruitment. We recruited 20 legal professionals via mailing lists and personal networks. Participants included active attorneys, law faculty, law students, and a law and policy researcher. Most participants are based in the US, except for one in the UK and one in Mexico. The cohort spans early-career to lawyers over 20 years of experience, with varying degrees of AI usage. Table 1 summarizes participants’ backgrounds and self-reported AI usage patterns. More detailed information is available at Appendix B.
@@
Construction of Cases. We manually sourced 33 cases from a combination of (1) the popular subreddit r/legaladvice (with wording edited slightly for anonymization and clarity), and (2) existing cases in legal practice familiar to our team member who is a practicing attorney. Our cases covered facets of law most relevant to lay users, spanning family law, criminal procedure, housing issues, and employment disputes. We selected cases that represented a diverse range of user intents (e.g., getting out of trouble, advocating for others, minimizing their costs), impacted third parties (e.g., employers, colleagues, landlords, family members, protesters), and degrees of damage (e.g., physical, financial, mental). This diversity was intended to elicit a wide range of discussion across legally and ethically sensitive contexts. Our cases can be viewed at https://github.com/Social-Futures-Lab/case-law-ai-policy/blob/main/data/cases.csv.
@@
Workshop Procedures. During the workshop, we presented 20 randomly-chosen cases along with 7 generic response strategies for AI responses on a shared Google document. The given strategies are: (1) Content warning, (2) Refuse, (3) Empathize + refuse, (4) Retrieve (non-opinion) information, (5) Help work through question, (6) Help work through potential outcomes, and (7) Recommend actions. We provided an example response for each strategy, and examples were derived from what we observed from OpenAI’s GPT-3.5 and 4. Both response strategies and corresponding examples are available in Appendix A. Because GPT tends to refuse to give detailed advice such as options (6) and (7), we drafted more specific answers complemented by known prompt engineering techniques such as drawing a hypothetical scenario (Greshake et al., 2023).
Figure 2 provides an overview of the collaborative document we gave the participants. After an introduction, each participant was given 10–15 minutes to freely choose 2–4 cases and (1) select the proper AI response strategies or produce their own preferred response and (2) the key dimensions impacting their decision in an individual workspace. Then, the experts had 30–35 minutes to discuss with each other why they chose certain response strategies and what dimensions they took into account to determine the proper strategies.
@@
Figure 2. Overview of case examples and LLM response strategies and examples provided to participants. Participants were given 10–15 minutes to review  20 legal case prompts on a shared document, select 2–4 cases to examine further, and specify appropriate LLM responses and influential considerations in their individual workspace on the same document.
@@
Analysis. We analyzed collaborative documents and transcripts using abductive coding (Tavory and Timmermans, 2014). Integrating both empirical data and available theory in an iterative process, our findings are informed by and enter into dialogue with literature from legal ethics (e.g., Haupt, 2019; Wendel, 2019) and ethical concerns in LLM interactions (e.g., Kim et al., 2023a; Valvoda et al., 2023). Our analysis synthesizes relevant aspects of these fields within the context of our research questions. Two authors initially coded 2 transcripts respectively and developed a codebook of dimensions and responses, informed by the Kim et al.’s human-AI-context framework (Kim et al., 2023b). The codebook was finalized through multiple all-author meetings. Following this, two coders independently analyzed the data and cross-checked each other’s work. In this process, both coders examined all documents and reached consensus on codes, rendering inter-rater reliability metrics unnecessary (McDonald et al., 2019).